Two angles are complementary. The measure of one angle is 21 more than twice the measure of the other angle. Find the measure of the angles.
Answer by hard(12)

To start off, complementary angles equal 90 degrees together.
In this problem:
Angle 2 = x
Angle 1 = 2x + 21 (because it's 2 times the measure of Angle 2 plus 21)
So -
x + 2x + 21 = 90
[Group like terms]
3x + 21 = 90
[Subtract 21 from both sides]
3x = 69
[Divide both sides by 3]
x = 23
So Angle 2 = 23 degrees.
Then, you "plug and chug".
Angle 1 = 2(23) + 21
Angle 1 = 46 + 21
Angle 1 = 67 degrees.
Check your answer to make sure it's correct.
67 + 23 = 90
Your answer is correct!